| 24 24 X-2 -=5 X+2 A) B) 9. 10} {=}
Answer #1

The given equation is

\\ \ \frac{24}{x-2}+\frac{24}{x+2}=5\\ \\ or, \frac{24(x+2)+24(x-2)}{(x-2)(x+2)}=5\\ \\or, \frac{24(2x)}{x^2-4}=5\\ \\ or, 48x=5(x^2-4)\\ \\ or, 5x^2-48x-20=0\\ \\ or, 5x^2-50x+2x-20=0\\ \\or, 5x(x-10)+2(x-10)=0\\ \\or, (5x+2)(x-10)=0\\ \\ either, (5x+2)=0, or, (x-10)=0\\ \\ \therefore \ x= -\frac{2}{5}, and\ \ x=10\\

\\ Ans :\\ \\ \bullet \ D) \ \ \left \{\ -\frac{2}{5}\ , \ 10 \ \right \}\\